OpenStudy (anonymous):

The longer leg of a right triangle is 8 more than the shorter leg. The hypotenuse is 8 more than the longer leg. Find all three sides.

OpenStudy (anonymous):

derp, help meh

OpenStudy (science0229):

You can use Pythagorean Theorem for this.

OpenStudy (anonymous):

derp

OpenStudy (science0229):

Put the shorter leg as x

OpenStudy (science0229):

Then the longer leg is x+8 and the hypotenuse is x+16

OpenStudy (anonymous):

hmm a^2 + b^2 = c

OpenStudy (science0229):

a^2 + b^2 = c^2

OpenStudy (science0229):

In this problem, a=x b=x+8 c=x+16

OpenStudy (science0229):

Solve that equation and find all 3 sides.

OpenStudy (anonymous):

oh wow thx dod

OpenStudy (anonymous):

it works

OpenStudy (science0229):

Warning: throw out the negative answer because length can't be negative.

OpenStudy (anonymous):

btw did you see, if you tag a lot of people @whpalmer4 @whydoihavetosignupl @EffImHighOffSkittles @eaileen @R00nnYBraiNsbiG @terenzreignz @uri @IloveYou* @primeralph @prettyboy20202 @ajprincess @AnElephant @SnuggieLad @snorl4x @Directrix @dan815 @divagirl421 @DarkKat @Dmitt @faix95 @GOODMAN @hobbs978 @jhonyy9 @Jack17 yu smartscore gets big

OpenStudy (science0229):

Also, shorter leg can't be longer than the longer leg.

OpenStudy (anonymous):

thx tho

OpenStudy (science0229):

welcome:)

OpenStudy (primeralph):